Opened 7 years ago

Closed 5 years ago

#456 closed flight processing (fixed)

EU09/06, flight day 242/2012, Porsmork

Reported by: knpa Owned by: tipo
Priority: alpha 4 medium Milestone: 2012 data processing completion
Component: Archiving Keywords:
Cc: Other processors:

Description (last modified by besm)

Data location: /users/rsg/arsf/arsf_data/2012/flight_data/greenland_iceland/EU09_06-2012_242_Porsmork

Data arrived from ARSF via SATA disk on 24/09/2012

EU09/06 Scientific objective:

EU09/06 Priority: a4m

PI: David Graham

Sensors:

Camera (20/12/12)
Eagle (20/12/12)
Hawk (20/12/12)
LiDAR (20/12/12)

Change History (19)

comment:1 Changed 7 years ago by knpa

  • Description modified (diff)
  • Summary changed from EU09/06, flight day 242a/2012, Porsmork to EU09/06, flight day 242/2012, Porsmork

comment:2 Changed 7 years ago by besm

Started nav processing, but .STA file didn't report the type of antenna used. After using teqc to convert data to rinex, it appears that the antenna is a LEIAX1202GG, with height of (apparently) 0.0m.

comment:3 Changed 7 years ago by besm

Basestation still wont process in grafnav. Processed navigation data using PPP instead. Still yet to see whether LiDAR will be accurate enough.

comment:4 Changed 7 years ago by besm

Pitch and roll values.

Pitch: -0.00312745
Roll: 0.000303727

comment:5 Changed 7 years ago by besm

LiDAR delivery created and ready to check.

comment:6 Changed 7 years ago by edfi

Starting LiDAR delivery check

comment:7 Changed 7 years ago by edfi

Check complete, all fine

comment:8 Changed 7 years ago by tipo

Started hyperspectral processing.

comment:9 Changed 7 years ago by tipo

Started RCD processing.

comment:10 Changed 7 years ago by tipo

RCD processing complete.

comment:11 Changed 7 years ago by tipo

SCT values:

LineEagleHawk
1-0.03-0.01
2-0.03-0.01
3-0.05-0.01
4-0.05-0.06
5-0.01-0.06
6-0.01-0.05
7-0.01-0.05
8-0.05-0.06
9-0.05-0.05
10-0.05-0.05
11-0.05-0.05

comment:12 Changed 7 years ago by tipo

Hyperspectral processing complete.

comment:13 Changed 7 years ago by edfi

Starting Hyperspectral delivery check

comment:14 Changed 7 years ago by edfi

Hyp check complete. Currently zipping mapped files. A some flightlines had areas of overflows, a note was made in the ReadMe

comment:15 Changed 7 years ago by besm

Sent to David Graham at Department of Geography, Loughborough University, Loughborough, Leics LE11 3TU. 20/12/12.

comment:16 Changed 7 years ago by besm

  • Description modified (diff)

comment:17 Changed 6 years ago by tipo

  • Component changed from Processing: general to Archiving
  • Owner set to tipo
  • Status changed from new to assigned

Beginning archiving

comment:18 Changed 6 years ago by knpa

NEODC currently ingesting.

comment:19 Changed 5 years ago by knpa

  • Resolution set to fixed
  • Status changed from assigned to closed

Archived by NEODC.

Closing ticket.

Note: See TracTickets for help on using tickets.